In the demand function Q= P^0.4 , calculate price elasticity of demand and identify type of good


1
Expert's answer
2021-05-04T07:35:02-0400

"Soln,"

Demand function "Q=P^{0.4}"


Price elasticity of demand"=\frac {p}{Q} \frac{dQ}{dP}"


"=\frac {dQ}{dP}=p^{0.4}"

"=\frac{dQ}{dP}=P^{-0.4-1}"


"E=\frac{P}{Q}\times\frac{dQ}{dP}"

"E=\frac{P}{P^{-0.4}}\times P^{-0.4-1}"


"=\frac{P}{P^{-0.4}}\times P^{-0.4}\times P^{-1}"

"=P\times P^{-1}=1"


"=PED=1"

This elasticity of demand is elastic, meaning as incomes increase there is also an increase in quantity demanded in the market. Hence it is a normal good.